OpenStudy (anonymous):

Is this statement true? log712 – log73 = log72 + log72

OpenStudy (dan815):

log laws

OpenStudy (dan815):

log(a)-Log(b) = Log(a/b)

OpenStudy (dan815):

and for addition you multiply

OpenStudy (dan815):

oh wait do u mean Logbase7

OpenStudy (dan815):

i was wondering why they wud have such an obviously false statment.. but now i see

ganeshie8 (ganeshie8):

\(\large \log_7 12 - \log_7 3 = \log_7 2 + \log_7 2\) like this maybe

OpenStudy (anonymous):

i think it is true

OpenStudy (anonymous):

log7(12/3) = log7 (2*2) with thew 7 i meant log base 7 log7(4)= log7 (4)